Momentum of a student with a cycle is 400kg m/s. Cycle is moving with a velocity of 5m/s. Find the mass of thestudent anda the cycle

Grade:9

2 Answers

Akshat Kumar tripathi
6 years ago
Let ,Mass of the man be M,And mass of the cycle be m,And we have the travelling velocity=5m/sNow (M+m)5 = 400 M+m =80 kg
Harsh
5 years ago
Let mass of the man is M and the mass of the cycle is m and the travelling velocity is 5m/sec. Now (M+m)5=400M+m=80KG